### v1

@@ 3 First application: image restoration
Images are often degraded during the data acquisition process. The degradation may involve blurring, information loss due to sampling, quantization effects, and various sources of noise. Image restoration, a popular application in computer vision, aims at recovering the original image from degraded data. UGMs are a popular tool for modeling the prior information such as smoothness in image restoration problems. In this framework, there is a node in the graph corresponding to each pixel in the image and the edges of the graph are chosen to enforce local smoothness conditions. The majority of the literature on solving the image restoration problem has focused on first-order UGMs, also known as, pairwise potentials; i.e., $|C|=2$ for all $C\in{\mathcal{C}}$. The most popular pairwise model for image restoration is the four-nearest neighbors model (see Figure 2 for an illustration) [43, 35, 40]. While it has been long recognized that higher-order UGMs are better suited for capturing properties of image priors, the complexity of solving Problem (B-UGM) has limited their use in practice. Almost all existing literature on higher-order binary UGMs tackles the MAP inference problem by reducing it to a binary quadratic optimization optimization to benefit from efficient optimization algorithms available for that problem class [1, 26, 38, 27, 28, 23].
@@ 3.1 Synthetic images
For each image type, we consider two sizes: small-size $15\times 15$ images and medium-size $100\times 100$ images. To generate blurred images, as in [8], we employ the bit-flipping noise model defined in Section 1 with $p\in[0.1:0.1:0.5]$. For each fixed $p$, we generate 50 random instances and report average relative optimality gaps and average CPU times. All LPs and IPs are solved with Gurobi 11.0 where all options are set to their default values. The relative optimality gaps and CPU times for different LP relaxations are shown in Figure 4 and Figure 5, respectively. In all instances, the clique LP returns a binary solution; i.e., Problem cliqueLP solves the original nonconvex Problem B-UGM. Therefore, for this test set, there is no need to run the multi-clique LP. As can be seen from Figure 4, the standard LP performs poorly in all cases. The flower LP leads to a moderate improvement in the quality of bounds, while the running LP significantly outperforms the flower LP. However, Figure 5 indicates that the computational cost of solving running LP is significantly higher than other LP relaxations. Hence, for this test set, clique LP is the best relaxation as it always solves the original problem and has the lowest computational cost among competitors.
@@
Figure 4: Relative optimality gap of LPs for synthetic images.
@@
To illustrate the importance of constructing strong LP relaxations for Problem B-UGM, let us comment on the computational cost of solving the IP; as we mentioned before the IP is obtained by adding binary requirements for all variables to Problem stdLP. We consider a $100\times 100$ ground truth image of type CEN and for each $p\in[0.1:0.1:0.5]$ we generate a blurred image. We then use Gurobi 11.0 to solve the IP and we set the time limit to 1800 seconds. Results are summarized in Table 2; none of the IP instances are solved to optimality within the time limit; we also report the relative gap of the IP upon termination. As can be seen from the table, in all cases the relative gap is larger than $200\%$. Interestingly, in all cases clique LP returns a binary solution in less than $5$ seconds.
@@ 3.2 QR codes
In this section, we demonstrate the effectiveness of the proposed LPs to restore an important type of real-world images: QR Codes. To this end, we utilize a more systematic approach for setting parameters of the inference problem so that we examine the quality of restored images. Henceforth, given a ground truth image and an algorithm for solving the image restoration problem, we measure the quality of the restored image in terms of partial recovery; i.e., the fraction of pixels that are identical in ground truth and restored images. Whenever an LP returns a fractional solution, we first round the solution to the closest binary point and then compute the partial recovery.
To capture the smoothness information for QR codes, we choose to learn the potential values from small-size and “mildly blurred” images. That is, we generate $10$ distinct $50\times 50$ QR codes and we use the bit-flipping noise model with $p=0.05$ to generate midly blurred instances. Next, we compute the average fraction of times $f_{i}$, $i\in[4]$ each group listed in Table 1 appears in these images. We then set $\varphi_{i}=-f_{i}$ for all $i\in[4]$. While this method is unrealistic for real-world problems as it assumes we have the ground truth image at hand, it imitates practical approaches in which practitioners consider a large database of somewhat clean images to learn the frequency of different potential patterns. Using more sophisticated techniques to learn potential values is beyond the scope of this paper.
@@
To test the performance of the proposed LP relaxations for restoring QR codes, we construct as the ground truth a $200\times 200$ QR code which contains a $100$-character-long text string. We set $p\in[0.1:0.1:0.5]$ and for each fixed $p$, we generate $50$ random instances. Results are shown in Figure 7. In addition to partial recovery rate, relative optimality gap, and CPU time, we also compare the tightness rate of different LPs; we define the tightness rate as the fraction of times each LP returns a binary solution. As can be seen from these figures, for this test set, standard LP and flower LP perform quite poorly, whereas, running LP, clique LP and multi-clique LP perform very well. Namely, the partial recovery rate of these three LPs is very close to that of the IP. Interestingly, multi-clique LP has the best tightness rate; however, in many instances for which running LP and clique LP return fractional solutions, the relative optimality gaps are very small, and the rounded binary solutions lead to similar partial recovery values to those of multi-clique LP. As before, the computational cost of solving running LP is significantly higher than other LPs. Hence for this test set multi-clique LP is the best option, followed closely by the clique LP.
@@ 4 Second application: decoding error-correcting codes
LDPC codes are often represented via UGMs; namely, each node of the graph corresponds to a message bit while each clique corresponds to a subset of bits with even parity. Gallager [24] introduced LDPC codes as error-correcting codes with three properties: $(i)$ all cliques have the same cardinality, denoted by $\beta$, $(ii)$ each node appears in the same number of cliques, denoted by $\gamma$, and $(iii)$ $\beta>\gamma$. Denoting by $n$ the number of message bits, an LDPC code is fully characterized by the triplet $(n,\beta,\gamma)$. Figure 9 illustrates an LDPC code.
@@
The above expression in turn can be equivalently written as:
@@ 4.1 The clique relaxation for decoding
The clique LP for Problem L-DCD is obtained by replacing the constraint $z\in S({\mathcal{G}})$ with $z\in\text{MP}^{\rm cl}({\mathcal{G}})$, where $\text{MP}^{\rm cl}({\mathcal{G}})$ is the clique relaxation and is defined in Section 2.4. In the following, we show that clique LP for Problem L-DCD has an interesting interpretation; namely, it is obtained by replacing the feasible region of Problem L-DCD corresponding to a single clique $C$ with its convex hull.
@@
To construct the convex hull of ${\mathcal{D}}(C)$ we make use of RLT as defined in [39]. That is, let $J_{1},J_{2}$ be any partition of $C$; define the factor $F(J_{1},J_{2})=\prod_{v\in J_{1}}{z_{v}}\prod_{v\in J_{2}}{(1-z_{v})}$. We first expand each $F(J_{1},J_{2})\geq 0$ and let $z_{e}=\prod_{v\in e}{z_{v}}$ for each product term to obtain linear inequalities (18). Subsequently, we multiply the equality constraint
@@
by each factor $F(J_{1},J_{2})$ and let $z_{e}=\prod_{v\in e}{z_{v}}$ to obtain a collection of linear equalities. Let us examine these equalities; two cases arise:
  • if $|J_{1}|$ is even, then multiplying $F(J_{1},J_{2})$ by equality (32) and using $z_{v}(1-z_{v})=0$ for any $v\in C$, we obtain the trivial equality $F(J_{1},J_{2})=F(J_{1},J_{2})$.
  • if $|J_{1}|$ is odd, then multiplying $F(J_{1},J_{2})$ by equality (32) and using $z_{v}(1-z_{v})=0$ for any $v\in C$, we obtain $F(J_{1},J_{2})=0$.
@@ 4.2 The parity polytope and the parity LP
In the following we show that clique LP is stronger than parity LP, in general. Denote by ${\mathcal{P}}({\mathcal{C}})$ the feasible region of Problem parLP. If ${\mathcal{C}}$ consists of a single clique $C$, then by Proposition 7:
@@
where ${\mathcal{H}}_{C}$ denotes the set of points satisfying equality (33). This implies that the feasible region of clique LP is contained in the feasible region of parity LP. As we detail next, this containment is often strict. We first examine the strength of clique LP. In [15], the authors generalize the decomposition result of Theorem 1 to account for additional constraints on multilinear sets:
Theorem 2 (Corollary 1 in [15]).
Let ${\mathcal{G}}$ be a hypergraph, and let ${\mathcal{G}}_{1}$,${\mathcal{G}}_{2}$ be section hypergraphs of ${\mathcal{G}}$ such that ${\mathcal{G}}_{1}\cup{\mathcal{G}}_{2}={\mathcal{G}}$ and ${\mathcal{G}}_{1}\cap{\mathcal{G}}_{2}$ is a complete hypergraph. Let $C({\mathcal{G}})$ be the set of points in $S({\mathcal{G}})$ that satisfy a number of constraints, each one containing only variables corresponding to nodes and edges only in ${\mathcal{G}}_{1}$ or only in ${\mathcal{G}}_{2}$. For $i=1,2$, let $C({\mathcal{G}}_{i})$ be the projection of $C({\mathcal{G}})$ in the space of $S({\mathcal{G}}_{i})$. Then, $C({\mathcal{G}})$ is decomposable into $C({\mathcal{G}}_{1})$ and $C({\mathcal{G}}_{2})$.
Thanks to Theorem 2 and Proposition 7, we can employ a similar line of arguments as in the proof of Proposition 4 to obtain a sufficient condition for sharpness of clique LP for decoding:
@@
Let ${\mathcal{G}}=\cup_{C\in{\mathcal{C}}}{{\mathcal{G}}_{C}}$ where ${\mathcal{G}}_{C}$ is a complete hypergraph with node set $C$. If ${\mathcal{C}}$ has the running intersection property, then
@@
If ${\mathcal{C}}=\{C_{1},C_{2}\}$, then ${\mathcal{C}}$ clearly has the running intersection property. However, as we show next, even in this case the parity relaxation ${\mathcal{P}}({\mathcal{C}})$ does not coincide with the convex hull.
@@ 4.3 Numerical Experiments
In this section, we compare the performance of different LP relaxations for decoding LDPC codes. We first describe how LDPC codes are generated [24]: an $(n,\beta,\gamma)$ LDPC code is often characterized by a parity-check matrix, an $m\times n$ binary matrix with $m:=\frac{n}{\gamma}\times\beta$, where each row contains $\beta$ ones and each column contains $\gamma$ ones. To construct a parity check matrix, we start by creating a matrix with all ones arranged in descending order; the $i$th row contains ones in columns $(i-1)\beta+1$ to $i\beta$. We then permute the columns of this matrix randomly and append it to the initial matrix. This permutation and appending is repeated $\gamma-1$ times to ensure each column contains $\gamma$ ones. The ones in each row of the matrix then correspond to the nodes of a clique in the UGM. It then follows that the UGM consists of $\frac{n}{\gamma}\times\beta$ cliques, each consisting of $\beta$ nodes. We assume an all-zero code as the ground truth code. As the first set of experiments, we consider a $(60,4,3)$ LDPC code. We use the bit-flipping noise with $p\in[0:0.01:0.2]$ and for each $p$ we generate $400$ random trials. We then compare the performance of different LPs with respect to tightness rate and partial recovery rate as defined before. For the multi-clique LP we set $m=4$; i.e., in our tests lifted odd-cycle inequalities (6) for cycles of cliques of length three and four are generated. Results are shown in Figure 10. As can be seen from this figure, standard LP and flower LP perform quite poorly, while running LP, clique LP, multi-clique LP, and parity LP perform well. Multi-clique LP is the best, followed by clique LP, followed by running LP, followed by the parity LP. As before the computational cost of solving running LP is significantly higher than other LPs. Motivated by these observations, in the next set of experiments we restrict our study to parity LP, clique LP, and multi-clique LP.
@@
We next consider three type of LDPC codes: $(120,4,3)$, $(120,5,4)$, and $(120,6,5)$. We set $p\in[0:0.01:0.2]$ and for each $p$ we generate $400$ random trials. Results are depicted in Figure 11. Overall, multi-clique LP and clique LP have better tightness rates than the parity LP, as the theory suggests. However, the differences, specially in terms of partially recovery rates become smaller as we increase the code length. This indeed, indicates the difficulty of solving this problem class; namely, by constructing stronger LP relaxations, the partial recovery rate of the decoder only marginally improves. These results also suggest that for a fixed code length, as we increase the clique size, the performance of all LPs degrade; while for a $(120,4,3)$ LDPC code all LPs manage to recover the ground truth with up to about $10\%$ corruption, for a $(120,6,5)$ LDPC code, this number decreases to about $5\%$ corruption.

### v2

@@ 3 First application: image restoration
Images are often degraded during the data acquisition process. The degradation may involve blurring, information loss due to sampling, quantization effects, and various sources of noise. Image restoration, a popular application in computer vision, aims at recovering the original image from degraded data. UGMs are a popular tool for modeling the prior information such as smoothness in image restoration problems. In this framework, there is a node in the graph corresponding to each pixel in the image and the edges of the graph are chosen to enforce local smoothness conditions. The majority of the literature on solving the image restoration problem has focused on first-order UGMs, also known as, pairwise models; i.e., $|C|=2$ for all $C\in{\mathcal{C}}$. The most popular pairwise model for image restoration is the four-nearest neighbors model (see Figure 2 for an illustration) [45, 36, 42]. While it has been long recognized that higher-order UGMs are better suited for capturing properties of image priors, the complexity of solving Problem (B-UGM) has limited their use in practice. Almost all existing literature on higher-order binary UGMs tackles the MAP inference problem by reducing it to a binary quadratic optimization optimization to benefit from efficient optimization algorithms available for that problem class [1, 27, 40, 28, 29, 23].
@@ 3.1 Synthetic images
For each image type, we consider two sizes: small-size $15\times 15$ images and medium-size $100\times 100$ images. To generate blurred images, as in [8], we employ the bit-flipping noise model defined in Section 1 with $p\in[0.1:0.1:0.5]$, where $[0.1:0.1:0.5]$ denotes a regularly-spaced vector between $0.1$ and $0.5$ using $0.1$ as the increment between elements. For each fixed $p$, we generate 50 random instances and report average relative optimality gaps and average CPU times. All LPs and IPs are solved with Gurobi 11.0, where all options are set to their default values. The relative optimality gaps and CPU times for different LP relaxations are shown in Figure 4 and Figure 5, respectively. In all instances, the clique LP returns a binary solution; i.e., Problem cliqueLP solves the original nonconvex Problem B-UGM. Therefore, for this test set, there is no need to run the multi-clique LP. As can be seen from Figure 4, the standard LP performs poorly in all cases. The flower LP leads to a moderate improvement in the quality of bounds, while the running LP significantly outperforms the flower LP. However, Figure 5 indicates that the computational cost of solving the running LP is significantly higher than other LP relaxations. Hence, for this test set, the clique LP is the best relaxation as it always solves the original problem and has the lowest computational cost among competitors.
@@
Figure 4: Relative optimality gap of LPs for synthetic images. The clique LP is not depicted in these pictures as this LP returns a binary solution in all cases, implying a zero optimality gap.
@@
To illustrate the importance of constructing strong LP relaxations for Problem B-UGM, let us comment on the computational cost of solving the IP; as we mentioned before, the IP is obtained by adding binary requirements for all variables to Problem stdLP. We consider a $100\times 100$ ground truth image of type CEN and for each $p\in[0.1:0.1:0.5]$ we generate a blurred image. We then use Gurobi 11.0 to solve the IP and we set the time limit to 1800 seconds. Results are summarized in Table 2; none of the IP instances are solved to optimality within the time limit; we also report the relative gap of the IP upon termination. As can be seen from the table, in all cases the relative gap is larger than $200\%$. Interestingly, in all cases the clique LP returns a binary solution in less than $5$ seconds.
Table 2: Performance of Gurobi 11.0 when solving the integer program for $100\times 100$ images of type CEN. The integer program is obtained by adding the integrality constraints to the standard LP.
@@ 3.2 QR codes
In this section, we demonstrate the effectiveness of the proposed LPs to restore an important type of real-world images: QR Codes. To this end, we utilize a more systematic approach for setting parameters of the inference problem so that we examine the quality of restored images. Henceforth, given a ground truth image and an algorithm for solving the image restoration problem, we measure the quality of the restored image in terms of partial recovery; i.e., the fraction of pixels that are identical in ground truth and restored images. Whenever an LP returns a fractional solution, we first round the solution to the closest binary point and then compute partial recovery.
To capture the smoothness information for QR codes, we choose to learn the potential values from small-size and “mildly blurred” images. That is, we generate $10$ distinct $50\times 50$ QR codes and we use the bit-flipping noise model with $p=0.05$ to generate mildly blurred instances. Next, we compute the average fraction of times $f_{i}$, $i\in[4]$ each group listed in Table 1 appears in these images. We then set $\varphi_{i}=-f_{i}$ for all $i\in[4]$. While this method is unrealistic for real-world problems as it assumes we have the ground truth image at hand, it imitates practical approaches in which practitioners consider a large database of somewhat clean images to learn the frequency of different potential patterns. Using more sophisticated techniques to learn potential values is beyond the scope of this paper.
@@
To test the performance of the proposed LP relaxations for restoring QR codes, we construct as the ground truth a $200\times 200$ QR code which contains a $100$-character-long text string. We set $p\in[0.1:0.1:0.5]$ and for each fixed $p$, we generate $50$ random instances. Results are shown in Figure 7. In addition to partial recovery rate, relative optimality gap, and CPU time, we also compare the tightness rate of different LPs; we define the tightness rate as the fraction of times each LP returns a binary solution. As can be seen from these figures, for this test set, the standard LP and the flower LP perform quite poorly, whereas, the running LP, the clique LP and the multi-clique LP perform very well. Namely, the partial recovery rate of these three LPs is very close to that of the IP. Interestingly, the multi-clique LP has the best tightness rate; however, in many instances for which the running LP and the clique LP return fractional solutions, the relative optimality gaps are very small, and the rounded binary solutions lead to similar partial recovery values to those of the multi-clique LP. As before, the computational cost of solving the running LP is significantly higher than other LPs. Hence for this test set the multi-clique LP is the best option, followed closely by the clique LP.
@@ 4 Second application: decoding error-correcting codes
LDPC codes are often represented via UGMs; namely, each node of the graph corresponds to a message bit while each clique corresponds to a subset of bits with even parity. Gallager [24] introduced LDPC codes as error-correcting codes with three properties: $(i)$ all cliques have the same cardinality, denoted by $\beta$, $(ii)$ each node appears in the same number of cliques, denoted by $\gamma$, and $(iii)$ $\beta>\gamma$. Denoting by $n$ the number of message bits, an LDPC code is fully characterized by the triplet $(n,\beta,\gamma)$. Figure 9 illustrates an LDPC code with $(n,\beta,\gamma)=(9,3,2)$.
@@
The clique potentials can be equivalently written as:
@@ 4.1 The clique relaxation for decoding
The clique LP for Problem L-DCD is obtained by replacing the constraint $z\in S({\mathcal{G}})$ with $z\in\text{MP}^{\rm cl}({\mathcal{G}})$, where $\text{MP}^{\rm cl}({\mathcal{G}})$ is the clique relaxation and is defined in Section 2.4. In the following, we show that the clique LP for Problem L-DCD has an interesting interpretation; namely, it is obtained by replacing the feasible region of Problem L-DCD corresponding to a single clique $C$ with its convex hull.
@@
To construct the convex hull of ${\mathcal{D}}(C)$, we make use of RLT as defined in [41]. That is, let $J_{1},J_{2}$ be any partition of $C$; define the factor $F(J_{1},J_{2})=\prod_{v\in J_{1}}{z_{v}}\prod_{v\in J_{2}}{(1-z_{v})}$. We first expand each $F(J_{1},J_{2})\geq 0$ and let $z_{e}=\prod_{v\in e}{z_{v}}$ for each product term to obtain linear inequalities (18). Subsequently, we multiply the equality constraint
@@
by each factor $F(J_{1},J_{2})$ and let $z_{e}=\prod_{v\in e}{z_{v}}$ to obtain a collection of equations. Let us examine these equations; two cases arise:
  • if $|J_{1}|$ is even, then multiplying $F(J_{1},J_{2})$ by equation (31) and using $z_{v}(1-z_{v})=0$ for any $v\in C$, we obtain the trivial equation $F(J_{1},J_{2})=F(J_{1},J_{2})$.
  • if $|J_{1}|$ is odd, then multiplying $F(J_{1},J_{2})$ by equation (31) and using $z_{v}(1-z_{v})=0$ for any $v\in C$, we obtain $F(J_{1},J_{2})=0$.
@@ 4.2 The parity polytope and the parity LP
In the following we show that the clique LP is stronger than the parity LP, in general. Denote by ${\mathcal{P}}({\mathcal{C}})$ the feasible region of Problem parLP. If ${\mathcal{C}}$ consists of a single clique $C$, then by Proposition 7:
@@
where ${\mathcal{H}}_{C}$ denotes the set of points satisfying equality (32). This implies that the feasible region of the clique LP is contained in the feasible region of the parity LP. As we detail next, this containment is often strict. We first examine the strength of the clique LP. In [15], the authors generalize the decomposition result of Theorem 1 to account for additional constraints on multilinear sets:
Theorem 2 ([15]).
Let ${\mathcal{G}}$ be a hypergraph, and let ${\mathcal{G}}_{1}$,${\mathcal{G}}_{2}$ be section hypergraphs of ${\mathcal{G}}$ such that ${\mathcal{G}}_{1}\cup{\mathcal{G}}_{2}={\mathcal{G}}$ and ${\mathcal{G}}_{1}\cap{\mathcal{G}}_{2}$ is a complete hypergraph. Let $C({\mathcal{G}})$ be the set of points in $S({\mathcal{G}})$ that satisfy a number of constraints, each one containing only variables corresponding to the nodes and edges only in ${\mathcal{G}}_{1}$ or only in ${\mathcal{G}}_{2}$. For $i=1,2$, let $C({\mathcal{G}}_{i})$ be the projection of $C({\mathcal{G}})$ in the space of $S({\mathcal{G}}_{i})$. Then, $C({\mathcal{G}})$ is decomposable into $C({\mathcal{G}}_{1})$ and $C({\mathcal{G}}_{2})$.
Thanks to Theorem 2 and Proposition 7, we can employ a similar line of arguments as in the proof of Proposition 4 to obtain a sufficient condition for sharpness of the clique LP for decoding:
@@
Denote by ${\mathcal{C}}$ the set of maximal cliques of a binary UGM. Let ${\mathcal{G}}=\cup_{C\in{\mathcal{C}}}{{\mathcal{G}}_{C}}$, where ${\mathcal{G}}_{C}$ is a complete hypergraph with node set $C$. If ${\mathcal{C}}$ has the running intersection property, then
@@
If ${\mathcal{C}}=\{C_{1},C_{2}\}$, then ${\mathcal{C}}$ clearly has the running intersection property, implying that the clique relaxation coincides with the convex hull. However, as we show next, even in this simple case, the parity relaxation ${\mathcal{P}}({\mathcal{C}})$ does not coincide with the convex hull.
@@ 4.3 Numerical Experiments
In this section, we compare the performance of different LP relaxations for decoding LDPC codes. We first describe how LDPC codes are generated [24]: an $(n,\beta,\gamma)$ LDPC code is often characterized by a parity-check matrix, an $m\times n$ binary matrix with $m:=\frac{n}{\gamma}\times\beta$, where each row contains $\beta$ ones and each column contains $\gamma$ ones. To construct a parity check matrix, we start by creating a matrix with all ones arranged in descending order; the $i$th row contains ones in columns $(i-1)\beta+1$ to $i\beta$. We then permute the columns of this matrix randomly and append it to the initial matrix. This permutation and appending is repeated $\gamma-1$ times to ensure each column contains $\gamma$ ones. The ones in each row of the matrix then correspond to the nodes of a clique in the UGM. It then follows that the UGM consists of $\frac{n}{\gamma}\times\beta$ cliques, each consisting of $\beta$ nodes. We assume an all-zero code as the ground truth code. As the first set of experiments, we consider a $(60,4,3)$ LDPC code. We use the bit-flipping noise with $p\in[0:0.01:0.2]$ and for each $p$ we generate $400$ random trials. We then compare the performance of different LPs with respect to tightness rate and partial recovery rate as defined before. For the multi-clique LP we set $m=4$; i.e., in our tests lifted odd-cycle inequalities (6) for cycles of cliques of length three and four are generated. Results are shown in Figure 10. As can be seen from this figure, the standard LP and the flower LP perform quite poorly, while the running LP, the clique LP, the multi-clique LP, and the parity LP perform well. The multi-clique LP is the best, followed by the clique LP, followed by the running LP, followed by the parity LP. As before the computational cost of solving the running LP is significantly higher than other LPs. Motivated by these observations, in the next set of experiments we restrict our study to the parity LP, the clique LP, and the multi-clique LP.
@@
We next consider three type of LDPC codes: $(120,4,3)$, $(120,5,4)$, and $(120,6,5)$. We set $p\in[0:0.01:0.2]$ and for each $p$ we generate $400$ random trials. Results are depicted in Figure 11. Overall, the multi-clique LP and the clique LP have better tightness rates than the parity LP, as the theory suggests. However, the differences, especially in terms of partial recovery rates become smaller as we increase the code length. This indeed, indicates the difficulty of solving this problem class; namely, by constructing stronger LP relaxations, the partial recovery rate of the decoder only marginally improves. These results also suggest that for a fixed code length, as we increase the clique size, the performance of all LPs degrade; while for a $(120,4,3)$ LDPC code all LPs manage to recover the ground truth with up to about $10\%$ corruption, for a $(120,6,5)$ LDPC code, this number decreases to about $5\%$ corruption.